OpenStudy (anonymous):

8 - (3+B) = B - 9 SHOW WORK PLEASE

OpenStudy (anonymous):

8-(3+b)=b-9 8-3-b=b-9 distributing the 1 that is before the ( ) -b -b 8-3-2b=-9 5-2b=-9 -5 -5 -2b=-14 B=7